The nth term of the series -9-6-3-......... and 16+14+12+....... are equal. Find the numbers of the terms in the series.

Grade:10

1 Answers

Rajat
4 years ago
For first AP, common difference= d = 3
nth term = -9 + (n-1)3
For second AP, common difference =d= -2
nth term =16+(n-1)(-2)
Equating both,
25=(n-1)(3+2)
Or, 5= (n-1) 
So, n= 6
Please approve